OpenStudy (anonymous):

Sand falls from a conveyer belt at the rate of 10 m^3/min onto the top of a conical pile. The height of the pile is always three-eights of the base diameter. How fast are the a) height and b) radius changing when the pile is 4 m high?

OpenStudy (anonymous):

I know that h = 3/4*r and dV/dt = 10 m^3/min, and V = pi r^2h

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):



zepdrix (zepdrix):

So they gave us a relationship between the height and the radius. That's very helpful, because we'll be able to Write the volume formula with ONLY H and V (to solve for h'), and then separately with ONLY R and V (to solve for r').

zepdrix (zepdrix):

\[\large V=\frac{ 1 }{ 3 }\pi r^2h=\frac{ 1 }{ 3 }\pi r^2(\frac{ 3 }{ 4 }r)\] See how we were able to write the RIGHT side in terms of only R?

OpenStudy (anonymous):

lol yea, so far I have dV/dt = \[(\frac{ 4 }{ 3 } h) h ^{2} * \frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

well i solved in terms of h to find dh/dt

zepdrix (zepdrix):

Oh you're doing the other one first :) k cool

zepdrix (zepdrix):

\[V=\frac{ 1 }{ 3 }\pi r^2h=\frac{ 1 }{ 3 }\pi (\frac{ 4 }{ 3 }h)^2h\] \[=\frac{ 1 }{ 3 }\pi (\frac{ 16 }{ 9 }h^3)\] Hmm I'm wondering if you plugged in your h correctly. Take a look at my setup.

OpenStudy (anonymous):

lol i haven't plugged in h yet

OpenStudy (anonymous):

oh wait 4/3 pi is squared?

OpenStudy (anonymous):

oops i see where i made my mistake lol